# K3 surfaces with real or complex multiplication

### v2

@@ 2. Integral and rational quadratic forms
(ii) The quadratic form $A_{2}\otimes_{{\mathbf{Z}}}{{\mathbf{Q}}}$ represents $-2$, hence it is isomorphic to $\langle-2,a\rangle$ for some $a\in\ QQ$. Since ${\rm det}(A_{2})=3$, we obtain $a=-6$. (iii) is proved by the same argument. The quadratic form $\langle-1,-1\rangle$ represents $-2$, hence it is isomorphic over ${\mathbf{Q}}$ to $\langle-2,a\rangle$ for some $a\in{\mathbf{Q}}$; since the determinant of $\langle-1,-1\rangle$ is 1, we have $a=-2$.
@@ 3. Some invariants
(i) ${\rm det}(H^{n})=(-1)^{n}$.
(ii) $w(H^{n})=0$ at $p$ if $p$ is a prime number $p\not=2$.
[…]
@@ 4. Hermitian forms and transfer
Let $(U,q)$ be a quadratic form over ${\mathbf{Q}}$, and suppose that $U$ has a structure of $E$-vector space. The following are equivalent
(i) For all $x,y\in U$ and all $\alpha\in E$, we have
@@
(ii) There exists a hermitian form $h:U\times U\to E$ such that for all $x,y\in U$, we have
@@ 5. Transfer
An algebraic number field $E$ is said to be a totally real field if for all embeddings $E\to\mathbf{C}$ we have $\sigma(E)\subset\mathbf{R}$; it is said to be a CM field if it is a totally imaginary field that is a quadratic extension of a totally real field.
@@
called the transfer of $W$, or more precisely of $(W,Q)$. Let $\Delta_{E}$ be the discriminant of $E$.
@@
Proof. (i) is clear, (ii) is proved in [vG 08], Lemma 4.5 (i) and part (iii) follows from [B 23], §19, determinant condition. ∎
@@
[…]
@@
Proof. See [B 23], §19, hyperbolicity condition. ∎
@@
Proof. If there exists a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$ then (i) clearly holds, property (ii) follows from Lemma 5.1 (ii), property (iii) from Lemma 5.4 and property (iv) from the above discussion. Conversely, suppose that conditions (i)-(iv) hold. The existence of a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$ follows from [B 23], Theorem 17.2. Indeed, since $E$ is a field, we have ${\mbox{{Sh}}}_{E}=0$. The hypotheses imply that condition (L 1) holds (see [B 23], Proposition 19.3). Therefore [B 23], Theorem 17.2 implies that there exists a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$. ∎
@@ 6. CM fields
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ .$
@@
Proof. Note that the signature of $V^{\prime}$ is $(1,1)$. Let $U$ be a quadratic form over ${\mathbf{Q}}$ such that $V\simeq U\oplus V^{\prime}$; this is possible by Lemma 1.2. The signature of $U$ is $(2,19)$. We have ${\rm det}(V^{\prime})=-\Delta_{E}^{\prime}$, hence ${\rm det}(U)=\Delta_{E}^{\prime}$. This implies that $U$ satisfies condition (ii) of Theorem 5.5; conditions (i) and (iv) clearly hold, so it remains to check condition (iii).
We have $w(V)=w(U)+w(V^{\prime})+({\rm det}(U),{\rm det}(V^{\prime}))$, and ${\rm det}(U)=\Delta_{E}^{\prime}$, ${\rm det}(V^{\prime})=-\Delta_{E}^{\prime}$, therefore $({\rm det}(U),{\rm det}(V^{\prime}))=(\Delta_{E},^{\prime}-\Delta_{E}^{\prime})=0$. This implies that $w(V)=w(U)+w(V^{\prime})$.
@@
If $m$ is even, then $\Delta_{E}$ is a square, hence $w(V^{\prime})=0$.
Suppose that $m$ is odd; since $md=20$, we have either $m=1$ and $d=20$ or $m=5$ and $d=4$, therefore $\Delta_{E}^{\prime}=\Delta_{E}$. If $p$ is a prime number such that $p\in S_{E}$, then $\Delta_{E}=1$ in ${\mathbf{Q}}_{p}^{\times}/{{\mathbf{Q}}_{p}^{\times}}^{2}$, hence $w(V^{\prime})=(a,\Delta_{E})=0$.
Hence in both cases $w(U)=w(V)=(-1,-1)$ at all $p\in S_{E}$, therefore $w(U)=w(H^{10})$ at $p$; hence $U$ satisfies condition (iii) of Theorem 5.5.
Therefore $U$ satisfies all the conditions of Theorem 5.5; hence there exists a hermitian form $W$ such that $U\simeq\mathrm{T}(W)$. ∎
@@
Let $V=V_{K3}$ and let $m$ be an integer with $m\geqslant 1$ such that $md\leqslant 20$. Then there exists a hermitian form $W$ over $E$ such that $\mathrm{T}(W)$ has signature $(2,18)$ and a quadratic form $V^{\prime}$ over ${\mathbf{Q}}$ such that
@@
$\bullet$ there exists a one-dimensional hermitian form $W$ over $E$ with $U\simeq\mathrm{T}(W)$;
$\bullet$ there exists a quadratic form $V^{\prime}$ over ${\mathbf{Q}}$ such that $U\oplus V^{\prime}\simeq V$.
@@
$\bullet$ If $v\in\Sigma$, then $v$ is inert in $E$,
$\bullet$ The cardinality of $\Sigma$ is even,
$\bullet$ If $v\in\Sigma$ is above the prime number $p$ and $v^{\prime}\in\Sigma$ with $v^{\prime}\not=v$ is above the prime number $p^{\prime}$, then $p\not=p^{\prime}$.
@@
There are infinitely many distinct sets of prime numbers $P(\Sigma)$ with $\Sigma$ as above.
@@
where the sum runs over the places $w$ of $E_{0}$ above $p$, the field $(E_{0})_{w}$ is the completion of $E_{0}$ at $w$, and ${\rm cor}_{(E_{0})_{w}/{{\mathbf{Q}}}_{p}}$ is the corestriction ${\rm Br}_{2}((E_{0})_{w})\to{\rm Br}_{2}({\mathbf{Q}}_{p})$ (see for instance [B 23], Proposition 12.4).
Let $\Sigma$ and $\Sigma^{\prime}$ be two sets as above with $P(\Sigma)\not=P(\Sigma^{\prime})$. Take $p\in P(\Sigma)$ with $p\not\in P(\Sigma^{\prime})$. By construction, we have $\underset{w|p}{\sum}{\rm cor}_{(E_{0})_{w}/{{\mathbf{Q}}}_{p}}(\lambda(\Sigma),\theta)\not=0$ and $\underset{w|p}{\sum}{\rm cor}_{(E_{0})_{w}/{{\mathbf{Q}}}_{p}}(\lambda(\Sigma^{\prime}),\theta)=0$, hence $w(U(\Sigma)\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p})\not=w(U(\Sigma^{\prime})\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p})$. This implies that $U(\Sigma^{\prime})$ and $U(\Sigma)$ are not isomorphic.
Since $d<20$, the existence of $V^{\prime}$ follows from Lemma 1.2.
@@
Let $k>0$ be an integer, and let $V=H^{3}\oplus\langle-2k\rangle$ and $m$ be an integer such that $md<6$, or $V=H^{3}\oplus I_{16}\oplus\langle-2k\rangle$ and $m$ be an integer such that $md<22$. Let $W$ be a hermitian form over $E$ of dimension $m$ such that ${\rm sign}(T(W))=(2,md-2)$. Then there exists a quadratic form over ${\bf Q}$ such that
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ .$
@@
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ .$
@@
We have ${\rm det}(V)=2k$ and $w(V)=(-1,-1)+(-1,-2k)=(-1,2k)$, and set $h=-2k\Delta_{E}$. Since $H$ is an orthogonal factor of $V$, there exists $x\in V$ such that $q(x,x)=h$; let $U$ be a quadratic form over $\bf Q$ such that $V\simeq U\oplus V^{\prime}$. We have ${\rm det}(U)={\rm det}(V){\rm det}(V^{\prime})=(2k)(-2k\Delta_{E})=-\Delta_{E}$; this implies that $w(V)=w(U)+(-\Delta_{E},-2k\Delta_{E})$. The signature of $U$ is $(2,md-2)$,
If $p\in S_{E}$, then $\Delta_{E}=1$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$, hence by the above computation we have $w(U)=(-1,2k)+(-1,-2k)=(-1,-1)$ at $p$; this implies that $w(U)=w(H^{3})$, respectively $w(H^{11})$. Therefore if $p\in S_{E}$, then $U\otimes_{\bf Q}{\bf Q}_{p}$ is isomorphic to $H^{3}\otimes_{{\mathbf{Q}}}{\mathbf{Q}}_{p}$, respectively $H^{11}\otimes_{{\mathbf{Q}}}{\mathbf{Q}}_{p}$. By Theorem 5.5 there exists a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$.
@@
Let $V=H^{3}\oplus\langle-2,-2\rangle$ and let $m$ be an integer with $m\geqslant 1$ such that $md\leqslant 6$. Then there exists a hermitian form $W$ over $E$ such that $\mathrm{T}(W)$ has signature $(2,md-2)$ and a quadratic form $V^{\prime}$ over ${\mathbf{Q}}$ such that
@@
Let $V=H^{3}\oplus I_{16}\oplus\langle-2,-6\rangle$, and let $m$ be an integer with $m\geqslant 1$ such that $md\leqslant 22$. Then there exists a hermitian form $W$ over $E$ such that $\mathrm{T}(W)$ has signature $(2,md-2)$ and a quadratic form $V^{\prime}$ over ${\mathbf{Q}}$ such that
$V\,\simeq\,{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ .$
Contrary to Proposition 6.8, Corollaries 6.9, 6.10 make no claim about the uniqueness of $V^{\prime}$, not even in the maximal dimensional case. Instead, we have the following:
@@
Let $V=H^{3}\oplus\langle-2,-2\rangle$ and let $m$ be an integer with $m\geqslant 1$ such that $md=6$. If $W$ is a hermitian form over $E$ such that $\mathrm{T}(W)$ has signature $(2,md-2)$ and if $V^{\prime}$ is a quadratic form over ${\mathbf{Q}}$ such that
$V\,\simeq\,{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ ,$
then $V^{\prime}\simeq\langle a,a\Delta_{E}\rangle$ with $a\in{\mathbf{Q}}^{\times}$ such that if $p\in S_{E}$, then $(-1,-a)=0$ at $p$.
@@
We have $w(V)=w(U)+w(V^{\prime})$, and $w(V)=0$, hence $w(V^{\prime})=w(U)$. If $p\in S_{E}$, then $w(U)=(-1,-1)$, therefore $w(V^{\prime})=w(\langle a,a\Delta_{E}\rangle)=(-1,-1)$. On the other hand, if $p\in S_{E}$, then $\Delta_{E}=1$, hence we obtain $w(V^{\prime})=(a,a)=(-1,-1)$, and this equivalent to $(-1,-a)=0$ for $p\in S_{E}$, as claimed. ∎
@@
Let $V=H^{3}\oplus I_{16}\oplus\langle-2,-6\rangle$, and let $m$ be an integer with $m\geqslant 1$ such that $md=22$. If $W$ is a hermitian form over $E$ such that $\mathrm{T}(W)$ has signature $(2,md-2)$ and if $V^{\prime}$ is a quadratic form over ${\mathbf{Q}}$ such that
$V\,\simeq\,{\mathrm{T}}(W)\oplus V^{\prime}\leavevmode\nobreak\ ,$
then $V^{\prime}\simeq\langle a,a\Delta_{E}\rangle$ with $a\in{\mathbf{Q}}^{\times}$ such that if $p\in S_{E}$, then $(-3,-2a)=0$ at $p$.
Proof. We have ${\rm det}(V)=-3$ and $w(V)=(2,6)$. Set $U={\mathrm{T}}(W)$; then ${\rm det}(U)=-\Delta_{E}$, therefore ${\rm det}(V^{\prime})=3\Delta_{E}$. Since ${\rm dim}(V^{\prime})=2$, this implies that $V^{\prime}\simeq\langle a,3a\Delta_{E}\rangle$ with $a\in{\mathbf{Q}}^{\times}$.
Let $p\in S_{E}$, then $\Delta_{E}=1$, and $V^{\prime}\simeq\langle a,3a\rangle$. Since $V\simeq U\oplus V^{\prime}$, we have $w(V)=w(U)+w(V^{\prime})+({\rm det}(U),{\rm det}(V^{\prime}))=(-1,-1)+w(V^{\prime})+(-1,3)=w(V^{\prime})+(-1,-3)$. We have $w(V)=(2,6)$, therefore we obtain $w(V^{\prime})=(2,6)+(-1,-3)=(-2,-6)$. Since $w(V^{\prime})=(a,3a)$, this is equivalent to $(a,3a)=(-2,-6)$, and this in turn can be reformulated as $(-3,-2a)=0$. ∎

### v3

@@ 3. Integral and rational quadratic forms
(ii) The quadratic form $A_{2}\otimes_{{\mathbf{Z}}}{{\mathbf{Q}}}$ represents 2-2, hence it is isomorphic to $\langle-2,a\rangle$ for some $a\in{\mathbf{Q}}$. Since ${\rm det}(A_{2})=3$, we obtain a=6a=-6. (iii) is proved by the same argument. The quadratic form $\langle-1,-1\rangle$ represents 2-2, hence it is isomorphic over $\mathbf{Q}$} to $\langle-2,a\rangle$ for some $a\in{\mathbf{Q}}$; since the determinant of $\langle-1,-1\rangle$ is 1, we have a=2a=-2.
@@ 4. Some invariants
  1. (1)
    ${\rm det}(H^{n})=(-1)^{n}$.
  2. (2)
    $w(H^{n})=0$ at $p$ if $p$ is a prime number $p\not=2$.
  3. (3)
    At $p=2$, we have $w(H^{n})=0$ if $n\equiv\ 0,1\ {\rm(mod\ 4)}$ and $w(H^{n})=1$ if $n\equiv\ 2,3\ {\rm(mod\ 4)}$.
@@ 5. Hermitian forms and transfer
Let $(U,q)$ be a quadratic form over $\mathbf{Q}$}, and suppose that $U$ has a structure of $E$-vector space. The following are equivalent:
  1. (1)
    For all $x,y\in U$ and all $\alpha\in E$, we have
@@
  2. (2)
    There exists a hermitian form $h:U\times U\to E$ such that for all $x,y\in U$, we have
@@ 6. Transfer
An algebraic number field $E$ is said to be a totally real field if for all embeddings $E\to\mathbf{C}$ we have $\sigma(E)\subset\mathbf{R}$; it is said to be a CM field if it is a totally imaginary field that is a quadratic extension of a totally real field. Recall that we introduced $\Delta_{E}$ to denote the discriminant of $E$.
@@
called the transfer of $W$, or more precisely of $(W,Q)$.
@@
Proof. (i) is clear, (ii) is proved in [vG 08], Lemma 4.5 (i) and part (iii) follows from [B 24a], §19, determinant condition. ∎
@@
In the above setting, $E\otimes_{{\mathbf{Q}}}{\mathbf{Q}}_{p}$ is called a split algebra. In the next section we will use that the discriminant of a split algebra is trivial.
@@
Let $E$ be a cyclotomic field, $E={\bf Q}(\zeta_{n})$. Then $S_{E}$ is the set of prime numbers $p$ such that the subgroup of $({\bf Z}/n{\bf Z})^{\times}$ generated by $p$ does not contain 1-1 (see [B 24b] Proposition 5.4, [B 24a] Proposition 31.2). If $n$ is a prime number with $n\equiv 3\ {\rm(mod\ 4)}$, then $p\in S_{E}$ if and only if $p$ is a square mod $n$ (see [B 24a], Corollary 31.3).
@@
Proof. See [B 24a], §19, hyperbolicity condition. ∎
@@
Proof. If there exists a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$ then (i) clearly holds, property (ii) follows from Lemma 6.1 (ii), property (iii) from Lemma 6.4 and property (iv) from the above discussion. Conversely, suppose that conditions (i)-(iv) hold. The existence of a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$ follows from [B 24a], Theorem 17.2. Indeed, since $E$ is a field, we have ${\mbox{{Sh}}}_{E}=0$. The hypotheses imply that condition (L 1) holds (see [B 24a], Proposition 19.3). Therefore [B 24a], Theorem 17.2 implies that there exists a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$. ∎
@@ 7. CM fields
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ .$
@@
Proof. Note that the signature of $V^{\prime}$ is $(1,1)$. Let $U$ be a quadratic form over $\mathbf{Q}$} such that $V\simeq U\oplus V^{\prime}$; this is possible by Lemma 2.3. The signature of $U$ is $(2,18)$. We have ${\rm det}(V^{\prime})=-\Delta_{E}^{\prime}$, hence ${\rm det}(U)=\Delta_{E}^{\prime}$. This implies that $U$ satisfies condition (ii) of Theorem 6.5; conditions (i) and (iv) clearly hold, so it remains to check condition (iii).
We have $w(V)=w(U)+w(V^{\prime})+({\rm det}(U),{\rm det}(V^{\prime}))$, and ${\rm det}(U)=\Delta_{E}^{\prime}$, ${\rm det}(V^{\prime})=-\Delta_{E}^{\prime}$, therefore $({\rm det}(U),{\rm det}(V^{\prime}))=(\Delta_{E}^{\prime},-\Delta_{E}^{\prime})=0$. This implies that $w(V)=w(U)+w(V^{\prime})$.
@@
If $m$ is even, then $\Delta_{E}^{\prime}$ is a square, hence $w(V^{\prime})=0$.
Suppose that $m$ is odd; since $md=20$, we have either $m=1$ and $d=20$ or $m=5$ and $d=4$, therefore $\Delta_{E}^{\prime}=\Delta_{E}$. If $p$ is a prime number such that $p\in S_{E}$, then $\Delta_{E}=1$ in ${\mathbf{Q}}_{p}^{\times}/{{\mathbf{Q}}_{p}^{\times}}^{2}$ by the comment just after Notation 6.2, hence $w(V^{\prime})=(a,\Delta_{E})=0$.
Hence in both cases w(U)=w(V)=(1,1)w(U)=w(V)=(-1,-1) at all $p\in S_{E}$ by Lemma 4.1. Here w(U)=(1,1)w(U)=(-1,-1), because $w(V)=w(U)+w(V^{\prime})+(\det(U),\det(V^{\prime}))$ and we already saw that this last term is trivial. Lemma 4.2 implies that $w(H^{10})=(-1,-1)$. Hence $w(U)=w(H^{10})$ at $p$. Since $U$ and $H^{10}$ have the same dimension and determinant, this implies that they are isomorphic. Therefore $U$ satisfies condition (iii) of Theorem 6.5.
In summary, $U$ satisfies all the conditions of Theorem 6.5; hence there exists a hermitian form $W$ such that $U\simeq\mathrm{T}(W)$. ∎
@@
Let $V=V_{K3}$ and let $m$ be an integer with $m\geqslant 1$ such that $md\leqslant 20$. Then there exists a hermitian form $W$ over $E$ such that $\mathrm{T}(W)$ has signature (2,md2)(2,md-2) and a quadratic form $V^{\prime}$ over $\mathbf{Q}$} such that
@@
  • $\bullet$
    there exists a one-dimensional hermitian form $W$ over $E$ with $U\simeq\mathrm{T}(W)$;
  • $\bullet$
    there exists a quadratic form $V^{\prime}$ over $\mathbf{Q}$} such that $U\oplus V^{\prime}\simeq V$.
@@
  • $\bullet$
    If $v\in\Sigma$, then $v$ is inert in $E$,
  • $\bullet$
    The cardinality of $\Sigma$ is even,
  • $\bullet$
    If $v\in\Sigma$ is above the prime number $p$ and $v^{\prime}\in\Sigma$ with $v^{\prime}\not=v$ is above the prime number $p^{\prime}$, then $p\not=p^{\prime}$.
@@
We shall use that there are infinitely many distinct sets of prime numbers $P(\Sigma)$ with $\Sigma$ as above. Indeed, since $E/E_{0}$ is a quadratic extension, there are infinitely many places $v$ of $E_{0}$ that are inert in $E$. It is easy to see that the two other conditions can also be satisfied, so that we obtain infinitely many different sets $\Sigma$, hence also $P(\Sigma)$ as stated.
@@
where the sum runs over the places $w$ of $E_{0}$ above $p$, the field $(E_{0})_{w}$ is the completion of $E_{0}$ at $w$, and ${\rm cor}_{(E_{0})_{w}/{{\mathbf{Q}}}_{p}}$ is the corestriction ${\rm Br}_{2}((E_{0})_{w})\to{\rm Br}_{2}({\mathbf{Q}}_{p})$ (see for instance [B 24a], Proposition 12.4).
Let $\Sigma$ and $\Sigma^{\prime}$ be two sets as above with $P(\Sigma)\not=P(\Sigma^{\prime})$. Take $p\in P(\Sigma)$ with $p\not\in P(\Sigma^{\prime})$. By construction, we have $\sum_{w\mid p}{\rm cor}_{(E_{0})_{w}/{{\mathbf{Q}}}_{p}}(\lambda(\Sigma),\theta)\not=0$ (since $(\lambda(\Sigma),\theta)=0$ at $v$ if and only if $v\not\in\Sigma$) and $\sum_{w\mid p}{\rm cor}_{(E_{0})_{w}/{{\mathbf{Q}}}_{p}}(\lambda(\Sigma^{\prime}),\theta)=0$, hence $w(U(\Sigma)\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p})\not=w(U(\Sigma^{\prime})\otimes_{{\mathbf{Q}}}{{\mathbf{Q}}}_{p})$. This implies that $U(\Sigma^{\prime})$ and $U(\Sigma)$ are not isomorphic. This concludes the proof of the proposition in case $d<20$.
@@
Let $k>0$ be an integer, and let $V=H^{3}\oplus\langle-2k\rangle$ and $m$ be an integer such that $md<6$, or $V=H^{3}\oplus I_{16}\oplus\langle-2k\rangle$ and $m$ be an integer such that $md<22$. Let $W$ be a hermitian form over $E$ of dimension $m$ such that ${\rm sign}(T(W))=(2,md-2)$. Then there exists a quadratic form $V^{\prime}$ over $\bf Q$} such that
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ .$
@@
$V\simeq{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ .$
@@
We have ${\rm det}(V)=2k$ and w(V)=(1,1)+(1,2k)=(1,2k)w(V)=(-1,-1)+(-1,-2k)=(-1,2k), and set $h=-2k\Delta_{E}$. Since $H$ is an orthogonal factor of $V$, there exists $x\in V$ such that $q(x,x)=h$; let $U$ be a quadratic form over $\bf Q$ such that $V\simeq U\oplus V^{\prime}$. We have ${\rm det}(U)={\rm det}(V){\rm det}(V^{\prime})=(2k)(-2k\Delta_{E})=-\Delta_{E}$; this implies that $w(V)=w(U)+(-\Delta_{E},-2k\Delta_{E})$. The signature of $U$ is (2,md2)(2,md-2),
If $p\in S_{E}$, then $\Delta_{E}=1$ in ${\mathbf{Q}}_{p}^{\times}/{\mathbf{Q}}_{p}^{\times 2}$, hence by the above computation we have w(U)=(1,2k)+(1,2k)=(1,1)w(U)=(-1,2k)+(-1,-2k)=(-1,-1) at $p$; this implies that $w(U)=w(H^{3})$, respectively $w(H^{11})$. Therefore if $p\in S_{E}$, then $U\otimes_{\bf Q}{\bf Q}_{p}$ is isomorphic to $H^{3}\otimes_{{\mathbf{Q}}}{\mathbf{Q}}_{p}$, respectively $H^{11}\otimes_{{\mathbf{Q}}}{\mathbf{Q}}_{p}$. By Theorem 6.5 there exists a hermitian form $W$ over $E$ such that $U\simeq\mathrm{T}(W)$.
@@
Let $V=H^{3}\oplus\langle-2,-2\rangle$ and let $m$ be an integer with $m\geqslant 1$ such that $md\leqslant 6$. Then there exists a hermitian form $W$ over $E$ such that $\mathrm{T}(W)$ has signature (2,md2)(2,md-2) and a quadratic form $V^{\prime}$ over $\mathbf{Q}$} such that
@@
Let $V=H^{3}\oplus I_{16}\oplus\langle-2,-6\rangle$, and let $m$ be an integer with $m\geqslant 1$ such that $md\leqslant 22$. Then there exists a hermitian form $W$ over $E$ such that $\mathrm{T}(W)$ has signature (2,md2)(2,md-2) and a quadratic form $V^{\prime}$ over $\mathbf{Q}$} such that
$V\,\simeq\,{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ .$
Contrary to Proposition 7.8, Corollaries 7.9, 7.10 make no claim about the uniqueness of $V^{\prime}$, not even in the maximal dimensional case. Instead, we have the following:
@@
Let $V=H^{3}\oplus\langle-2,-2\rangle$ and let $m$ be an integer with $m\geqslant 1$ such that $md=6$. If $W$ is a hermitian form over $E$ such that $\mathrm{T}(W)$ has signature (2,md2)(2,md-2) and if $V^{\prime}$ is a quadratic form over $\mathbf{Q}$} such that
$V\,\simeq\,{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ ,$
then $V^{\prime}\simeq\langle a,a\Delta_{E}\rangle$ with $a\in{\mathbf{Q}}^{\times}$ such that if $p\in S_{E}$, then (1,a)=0(-1,-a)=0 at $p$.
@@
We have $w(V)=w(U)+w(V^{\prime})$, and $w(V)=0$, hence $w(V^{\prime})=w(U)$. If $p\in S_{E}$, then w(U)=(1,1)w(U)=(-1,-1) locally at $p$, therefore $w(V^{\prime})=w(\langle a,a\Delta_{E}\rangle)=(-1,-1)$. On the other hand, if $p\in S_{E}$, then $\Delta_{E}=1$ locally at $p$ by the comment after Notation 6.2; hence we obtain $w(V^{\prime})=(a,a)=(-1,-1)$, and this equivalent to (1,a)=0(-1,-a)=0 for $p\in S_{E}$, as claimed (cf. Example 1.3 (1)). ∎
@@
Let $V=H^{3}\oplus I_{16}\oplus\langle-2,-6\rangle$, and let $m$ be an integer with $m\geqslant 1$ such that $md=22$. If $W$ is a hermitian form over $E$ such that $\mathrm{T}(W)$ has signature (2,md2)(2,md-2) and if $V^{\prime}$ is a quadratic form over $\mathbf{Q}$} such that
$V\,\simeq\,{\mathrm{T}}(W)\oplus V^{\prime}\penalty 10000\ ,$
then $V^{\prime}\simeq\langle a,a\Delta_{E}\rangle$ with $a\in{\mathbf{Q}}^{\times}$ such that if $p\in S_{E}$, then (3,2a)=0(-3,-2a)=0 at $p$.
Proof. We have ${\rm det}(V)=-3$ and $w(V)=(2,6)$. Set $U={\mathrm{T}}(W)$; then ${\rm det}(U)=-\Delta_{E}$ because $m$ and $d_{0}$ are both odd, therefore ${\rm det}(V^{\prime})=3\Delta_{E}$. Since ${\rm dim}(V^{\prime})=2$, this implies that $V^{\prime}\simeq\langle a,3a\Delta_{E}\rangle$ with $a\in{\mathbf{Q}}^{\times}$.
Let $p\in S_{E}$, then $\Delta_{E}=1$, and $V^{\prime}\simeq\langle a,3a\rangle$. Since $V\simeq U\oplus V^{\prime}$, we have $w(V)=w(U)+w(V^{\prime})+({\rm det}(U),{\rm det}(V^{\prime}))=(-1,-1)+w(V^{\prime})+(-1,3)=w(V^{\prime})+(-1,-3)$. We have $w(V)=(2,6)$, therefore we obtain $w(V^{\prime})=(2,6)+(-1,-3)=(-2,-6)$. Since $w(V^{\prime})=(a,3a)$, this is equivalent to (a,3a)=(2,6)(a,3a)=(-2,-6), and this in turn can be reformulated as (3,2a)=0(-3,-2a)=0 (cf. Example 1.3 (2)). ∎